# LiDAR-GS: Real-time LiDAR Re-Simulation using Gaussian Splatting

### v2

@@ Abstract
We present LiDAR-GS, a Gaussian Splatting (GS) method for real-time, high-fidelity re-simulation of LiDAR scans in public urban road scenes. Recent GS methods proposed for cameras have achieved significant advancements in real-time rendering beyond Neural Radiance Fields (NeRF). However, applying GS representation to LiDAR, an active 3D sensor type, poses several challenges that must be addressed to preserve high accuracy and unique characteristics. Specifically, LiDAR-GS designs a differentiable laser beam splatting, using range-view representation for precise surface splatting by projecting lasers onto micro cross-sections, effectively eliminating artifacts associated with local affine approximations. Furthermore, LiDAR-GS leverages Neural Gaussian Representation, which further integrate view-dependent clues, to represent key LiDAR properties that are influenced by the incident direction and external factors. Combining these practices with some essential adaptations, e.g., dynamic instances decomposition, LiDAR-GS succeeds in simultaneously re-simulating depth, intensity, and ray-drop channels, achieving state-of-the-art results in both rendering frame rate and quality on publically available large scene datasets when compared with the methods using explicit mesh or implicit NeRF. Our source code is publicly available at https://www.github.com/cqf7419/LiDAR-GS.
I Introduction
Closed-loop simulation [1, 2] is the basis for verifying sensor-to-control autonomous driving systems [3, 4]. In recent advances, camera re-simulation capabilities based on Gaussian Splatting [5] have gradually been integrated into simulation platforms [6], enabling feasible and applicable real-time rendering in such reconstructed scenes. However, since L4 autonomy still depends on LiDAR to ensure safety in challenging driving conditions, there is a high demand for supporting LiDAR re-simulation to verify multi-modal perception algorithms [7, 8] and end-to-end algorithms [3].
@@ I Introduction
Figure 1: We present LiDAR-GS, a LiDAR Gaussian Splatting for explicitly representing scenes and re-simulating LiDAR frames in real-time.
Our approach addresses these four differences to bring-up LiDAR-GS accordingly: (1) We choose range-view representation [10] to organize laser beams and design a differentiable laser beam splatting method, projecting 3D Gaussians onto beam-aligned micro-planes instead of the unified NDC. This eliminates affine approximation errors while preserving geometric fidelity. (2) For high-fidelity modeling and rendering on these LiDAR properties sensitive to the incident direction and distance, we propose to use Neural Gaussian Representation instead of the Spherical Harmonics, reaching higher quality of re-simulation according to our experiments (Table V and Figure 9). (3) For simulating the ray-drop characteristics, we assign a learnable ray-drop probability to each Gaussian primitive, letting forward rendering to judge the validity of emitted laser beam. (4) For vertical sparsity, we introduce spatial scale regularization and compact bounding-box strategies to constrain Gaussian scales and pixel coverage, ensuring precise gradient propagation. A dynamic decomposition mechanism further adapts the model to real-world driving scenarios without compromising real-time performance.
@@
II Related Work
Among applicable simulators [2], CARLA [13] builds the traditional render engine upon manufactured virtual scenes with assets modeling, texture editing, and shading parameters tuning. Such a high expense hinders a large-scale generation of virtual scenes. Latest advancements such as MARS [14] and UniSim [1] take advantage of NeRF’s realistic rendering ability, enabling re-simulating sequences captured from real-world environments and guaranteeing a smaller sim-to-real gap [15, 16]. However, these NeRF-based simulators are comparably less efficient than GS-based rendering, requiring more effort to bring-up a real-time simulation system, and they neither support structuralized scene editing due to their implicit representation. Consequently, several approaches substitute the scene representation to GS for cameras (Section II-A), whereas such a representation has not been thoroughly considered for LiDAR (Section II-B).
II-A Gaussian Splatting for Cameras
@@
II-B LiDAR Re-simulation Methods
@@
NeRF based methods. Recent LiDAR re-simulation methods [30, 31, 32, 33, 34] extend the NeRF formulation for novel LiDAR view synthesis. Among them, LiDAR-NeRF [30] and NFL [32] firstly propose the differentiable LiDAR re-simulation framework, which treats the oriented LiDAR laser beams as a set of rays and render 3D points and intensities in a similar manner as RGB. DyNFL [31] and LiDAR4D [34] extend to dynamic scenes. AlignMiF [33] further proposes to incorporate and align multimodal inputs for better LiDAR re-simulation. Through our experiments, we found refactoring LiDAR re-simulation with Gaussian Splatting follows the trend of camera re-simulation discussed in section II-A.
Concurrent GS approaches. Recent concurrent studies, GS-LiDAR [35] and LiDAR-RT [36], also explored LiDAR data synthesis. Different from them, our work adopts Neural Gaussian Representation and Laser beam Splatting to model the scene and achieves more comprehensive verification on re-simulation, novel view synthesis, beam changing and vehicle place-in.
III Preliminaries
Figure 2: The overview of LiDAR-GS pipeline. We choose to represent LiDAR frames as range-view images (Section III-A), and use Gaussian Splatting to perform both reconstruction and rendering (Section III-B). Specifically, we choose to use Neural Gaussian Representation (NGR) with four MLPs to learn final attributes for each Gaussian (Section IV-A), where projecting and rasterizing these LiDAR Gaussians are performed through our proposed Laser Beam Splatting strategy (Section IV-B). Dynamic instances are separately reconstructed in their stand-alone coordinates for final stitching (Section IV-C).
III-A Range-view For LiDAR Representation
@@
IV LiDAR Gaussian Splatting
IV-A LiDAR-GS Overview
@@
where we found MLP achieves better deduction than Spherical Harmonics when an input direction $\mathbf{d}^{\prime}$ and the distance-of-flight $d$ is given (See Table V).
@@
IV-B Laser Beam Splatting
@@
Since there is no longer a normalized image plane or sphere to determine the range of affected pixels for each Gaussian during projection, we employed a compact Axis-Aligned Bounding-Box (AABB) to ascertain the pixel coverage of each Gaussian. Specifically, we establish the major axes $\mathbf{a}$ and minor axes $\mathbf{b}$ of the Gaussian projection on the projection plane and transform it into a rectangular bounding box $(b_{h},b_{w})$, where $b_{h}=\mathbf{b}/\tan\phi,b_{w}=\mathbf{a}/\tan\theta$. Due to the sparsity of the laser beam in the vertical FOV, the vertical angle $\phi$ is usually bigger than the horizontal angle $\theta$, leading to significant variations in gradient. Therefore, compared to directly using the maximum radius as the selection box [5], our strategy reduces the unnecessary pixel coverage, which is beneficial for optimizing the vertical direction. The ablation study in Table V confirms the effectiveness of this strategy.
@@
5:     $d\leftarrow|\mathbf{x}-\mathbf{o}|,\ \ \mathbf{d}^{\prime}\leftarrow(\mathbf{x}-\mathbf{o})/d$triangleright Depth and Ray
6:     $n_{1}\leftarrow n_{1}\perp\mathbf{d}^{\prime},\ \ n_{2}\leftarrow n_{2}\perp\mathbf{d}^{\prime},n_{1}$triangleright Basis of plane
7:     $\Pi,\mathbf{J}\leftarrow n_{1},n_{2}$triangleright Tangent plane and Transformation
8:     ${\Sigma}^{\prime}\leftarrow\mathbf{J}\mathbf{W}{\Sigma}\mathbf{W}^{\top}\mathbf{J}^{\top}$triangleright Projection of covariance
9:     $b_{h},b_{w}\leftarrow{\Sigma}^{\prime}$triangleright Range of affected pixels
@@
13:     $\mathbf{N}_{\mathbf{d}}\leftarrow\xi:\{b_{h},b_{w}\}$triangleright Depth buffering
14:     $\mathbf{p}:[\bar{\rho},\bar{d},\bar{r}]\leftarrow[0,0,0]$triangleright Render Target
@@
17:        $\partial\mathbf{x}\leftarrow d,\epsilon$triangleright Back-projection
18:        $\alpha_{i}\leftarrow\alpha_{\xi_{i}}\cdot\mathbb{G}_{i}(\partial\mathbf{x})$triangleright Splatting and $\alpha$-weighting
19:        $[\bar{\rho},\bar{d},\bar{r}]\leftarrow\alpha_{i}\sum_{j=1}^{i-1}(1-\alpha_{j})[\rho_{\xi_{i}},d_{\xi_{i}},r_{\xi_{i}}]$
@@
22:        $\bar{d},\bar{\rho}\leftarrow 0$triangleright Ray-drop
@@
IV-C Dynamic Instances Decomposition
@@
V Experiments
V-A Experimental Setting
Real-world Large Scene Datasets. We verified the LiDAR-GS using 12 sequences from Waymo Open Dataset [11] and KITTI-360 Dataset [12], which are public real-world datasets for autonomous driving scenarios. Consistent with previous LiDAR re-simulation methods [30, 31], we utilized 50 continuous frames per sequence, designating every 10th frame for validation and the remaining frames for training. Please refer to Table I for the specific sequences chosen.
TABLE I: Sequences from Waymo and KITTI-360 for evaluation.
@@ V-A Experimental Setting
Baseline and Implementation. For static dataset, we used the official settings to conduct experiments with LiDAR-NeRF [30] and AlignMiF [33]. For NFL [32] which have not yet released their source-code, we stripped the dynamic graph from DyNFL [31] and followed its settings to reproduce the results. Additionally, we maintained the experimental settings as described in LiDAR-NeRF [30] to reproduce the results on PCGen [29] and LiDARsim [28]. For dynamics dataset, we reproduced the results using the same data sequences and official settings as DyNFL [31] and LiDAR4D [34], while also recording the official results of LiDARsim [28] and UniSim [1] from DyNFL [31].
@@
TABLE II: Average quantitative results for static scenes from the Waymo Open Dataset and KITTI-360.
@@
V-B Main Results
Figure 7: For dynamic scenes, qualitative comparison between LiDAR-GS and other methods on Waymo sequence.
TABLE III: Average quantitative results for dynamic scenes.
@@ V-B Main Results
TABLE IV: Average training and inference costs. We reach real-time rendering for LiDAR re-simulation tasks.
Method Train (min)$\downarrow$ Inference (fps)$\uparrow$
@@
Table II and Table III present our quantitative results on the public static dataset and dynamic dataset, respectively. As shown, our proposed method LiDAR-GS achieves best quantitative performance among most metrics. Meanwhile, Table IV demonstrates that our approach significantly reduces training and inference time compared to existing explicit reconstruction and NeRF methods. This improvement marks a milestone that LiDAR re-simulation can now achiev the same or even higher frame rate than real-world sensors, typically at 10Hz, making large-scale re-simulation of real-world sequences feasible.
Results on Static Dataset. The quantitative comparison on the Waymo dataset with the static scene is displayed in Table II, illustrating LiDAR-GS has achieved state-of-the-art in most indicators. As shown in Figure 6, LiDAR-GS is superior to previous methods in modeling scene surfaces and intensity, such as roads, vehicles, and poles. When testing on KITTI-360, we found that LiDAR-GS performs slightly worse on the CD metric compared to LiDAR-NeRF and AlignMiF. This performance drop may be attributed to sparse initialization in occluded or distant areas, which is not friendly to the optimization. And NeRF-based method can always maintain a continuous scene representation.
Results on Dynamic Dataset. The quantitative comparison on the Waymo dataset with dynamic objects is displayed in Table III. We can find our proposed method achieves remarkable performance on most evaluation metrics, underscoring the realism of our re-simulated LiDAR scans. Although LiDAR-GS scores lower than LiDAR4D on some intensity metrics (attributed to its two-stage convolutional network optimization), our rendered results, as shown in Figure 7, appear more realistic.
Results on Novel View Synthesis. We adjusted the free-viewpoint by lateral shifting (Y) and elevating (Z), or change the number of laser beams (from 64 to 32) for LiDAR re-simulation on novel views as shown in Figure 8, demonstrating our effectiveness of NVS in driving scenarios.
V-C Ablation Results
TABLE V: Ablations for the proposed method on Waymo1067.
@@ V-C Ablation Results
Ablations on Neural Gaussian Representation. We conducted an ablation experiment to remove the NGR. Table V supports our assertions in Section IV-A. The introduction of NGR significantly enhances the capability of scene representation. It can effectively consider the impact of incident direction and distance. As shown in Figure 9, if Gaussian properties are not related to above factors, it cannot identify positions of ray-drop, leading to incorrect optimization of intensity and depth, such as road blind spots.
@@
Ablations on Laser Beam Splatting. We conducted an ablation experiment to remove the LBS. Specifically, we referred to the camera projection method of the vanilla 3DGS [5] and projected Gaussian onto unified pano projection planes for rasterization. Table V demonstrates that LBS is more suitable for LiDAR applications. The improvements provided by LBS are illustrated in Figure 9, confirming our assertions in LABEL:{sec:method:laser-beam-splatting}. Additionally, we implemented LBS using CUDA, resulting in only slight reduction in efficiency. The costs of projections are detailed in Table IV.
Ablations on AABB. The results in Table V demonstrate that avioding unnecessary pixel coverage by using AABB for Gaussian projection is beneficial for further convergence.
@@
Figure 10: Ablations on the necessity of adding ray-drop estimation.
@@
Discussion on the gap between real and simulation. To gauge the gap between real and simulated scans, we employed detection model, INT [41], on the same trajectory of dynamic data to evaluate the difference in detection results between two point cloud detection results relative to manually annotated GT. Table VI presents that this difference is negligible in this task, where mAP(mean Average Precision), mF1(mean F1 Score) and mAVE(mean Average Velocity Error) are used as detection metric.
TABLE VI: Objects detection using point clouds.
@@
VI Conclusion
In this paper, we present the first LiDAR Gaussian Splatting method for real-time high-fidelity re-simulation of sensor scans in real-world urban road scenes. LiDAR-GS choose range-view representation to organize laser beams, and innovatively proposes a differentiable laser beam splatting based on micro cross-section plane. Furthermore, LiDAR-GS precisely simulates LiDAR properties using Neural Gaussian Representation and introducing some strategies dedicated to Gaussian laser beam, including AABB on projection constraints, and spatial scale regularization. LiDAR-GS has been verified in multiple dynamic and static scenes, and performs well. In summary, this GS-based rendering LiDAR re-simulation method will accelerate the progress of autonomous driving in data closure.

### v3

@@ Abstract
We present LiDAR-GS, a Gaussian Splatting (GS) method for real-time, high-fidelity re-simulation of LiDAR scans in public urban road scenes. Recent GS methods proposed for cameras have achieved significant advancements in real-time rendering beyond Neural Radiance Fields (NeRF). However, applying GS representation to LiDAR, an active 3D sensor type, poses several challenges that must be addressed to preserve high accuracy and unique characteristics. Specifically, LiDAR-GS designs a differentiable laser beam splatting, using range-view representation for precise surface splatting by projecting lasers onto micro cross-sections, effectively eliminating artifacts associated with local affine approximations. Furthermore, LiDAR-GS leverages Neural Gaussian Representation, which further integrate view-dependent clues, to represent key LiDAR properties that are influenced by the incident direction and external factors. Combining these practices with some essential adaptations, e.g., dynamic instances decomposition, LiDAR-GS succeeds in simultaneously re-simulating depth, intensity, and ray-drop channels, achieving state-of-the-art results in both rendering frame rate and quality on publically available large scene datasets when compared with the methods using explicit mesh or implicit NeRF. Our source code is publicly available.
Figure 1: We present LiDAR-GS, the LiDAR Gaussian Splatting for re-simulating LiDAR frames rendered in real-time.
Keywords: lidar reconstruction and re-simulation, rendering, gaussian splatting.
1 Introduction
@@
Our approach addresses these four differences to bring-up LiDAR-GS accordingly: (1) We choose range-view representation [11] to organize laser beams and design a differentiable laser beam splatting method, projecting 3D Gaussians onto beam-aligned micro-planes instead of the unified NDC. This eliminates affine approximation errors while preserving geometric fidelity. (2) For high-fidelity modeling and rendering on these LiDAR properties sensitive to the incident direction and distance, we propose to use Neural Gaussian Representation instead of the Spherical Harmonics, reaching higher quality of re-simulation according to our experiments (Table 5 and Figure 8). (3) For simulating the ray-drop characteristics, we assign a learnable ray-drop probability to each Gaussian primitive, letting forward rendering to judge the validity of emitted laser beam. (4) For vertical sparsity, we introduce spatial scale regularization and compact bounding-box strategies to constrain Gaussian scales and pixel coverage, ensuring precise gradient propagation. A dynamic decomposition mechanism further adapts the model to real-world driving scenarios without compromising real-time performance.
@@
2 Related Work
Among applicable simulators [14], CARLA [3] builds the traditional render engine upon manufactured virtual scenes with assets modeling, texture editing, and shading parameters tuning. Such a high expense hinders a large-scale generation of virtual scenes. Latest advancements such as MARS [30] and UniSim [32] take advantage of NeRF’s realistic rendering ability, enabling re-simulating sequences captured from real-world environments and guaranteeing a smaller sim-to-real gap [19, 16]. However, these NeRF-based simulators are comparably less efficient than GS-based rendering, requiring more effort to bring-up a real-time simulation system, and they neither support structuralized scene editing due to their implicit representation. Consequently, several approaches substitute the scene representation to GS for cameras (Section 2.1), whereas such a representation has not been thoroughly considered for LiDAR (Section 2.2).
2.1 Gaussian Splatting for Cameras
@@
2.2 LiDAR Re-simulation Methods
@@
NeRF based methods. Recent LiDAR re-simulation methods [26, 29, 7, 27, 34] extend the NeRF formulation for novel LiDAR view synthesis. Among them, LiDAR-NeRF [26] and NFL [7] firstly propose the differentiable LiDAR re-simulation framework, which treats the oriented LiDAR laser beams as a set of rays and render 3D points and intensities in a similar manner as RGB. DyNFL [29] and LiDAR4D [34] extend to dynamic scenes. AlignMiF [27] further proposes to incorporate and align multimodal inputs for better LiDAR re-simulation. Through our experiments, we found refactoring LiDAR re-simulation with Gaussian Splatting follows the trend of camera re-simulation discussed in section 2.1.
Contemporaneous methods. Contemporaneous publicly available work includes GS-LiDAR [8] and LiDAR-RT [35]. On the same sequences, our method achieves a reconstruction efficiency that is $7.1\times$ faster.
3 Preliminaries
Figure 2: The overview of LiDAR-GS pipeline. We choose to represent LiDAR frames as range-view images (Section 3.1), and use Gaussian Splatting to perform both reconstruction and rendering (Section 3.2). Specifically, we choose to use Neural Gaussian Representation (NGR) with four MLPs to learn final attributes for each Gaussian (Section 4.1), where projecting and rasterizing these LiDAR Gaussians are performed through our proposed Laser Beam Splatting strategy (Section 4.2). Dynamic instances are separately reconstructed in their stand-alone coordinates for final stitching (Section 4.3).
3.1 Range-view For LiDAR Representation
@@
4 LiDAR Gaussian Splatting
4.1 LiDAR-GS Overview
@@
where we found MLP achieves better deduction than Spherical Harmonics when an input direction $\mathbf{d}^{\prime}$ and the distance-of-flight $d$ is given (See Table 5).
@@
4.2 Laser Beam Splatting
@@
Since there is no longer a normalized image plane or sphere to determine the range of affected pixels for each Gaussian during projection, we employed a compact Axis-Aligned Bounding-Box (AABB) to ascertain the pixel coverage of each Gaussian. Specifically, we establish the major axes $\mathbf{a}$ and minor axes $\mathbf{b}$ of the Gaussian projection on the projection plane and transform it into a rectangular bounding box $(b_{h},b_{w})$, where $b_{h}=\mathbf{b}/\tan\phi,b_{w}=\mathbf{a}/\tan\theta$. Due to the sparsity of the laser beam in the vertical FOV, the vertical angle $\phi$ is usually bigger than the horizontal angle $\theta$, leading to significant variations in gradient. Therefore, compared to directly using the maximum radius as the selection box [10], our strategy reduces the unnecessary pixel coverage, which is beneficial for optimizing the vertical direction. The ablation study in Table 5 confirms the effectiveness of this strategy.
@@
5:  $d\leftarrow|\mathbf{x}-\mathbf{o}|,\ \ \mathbf{d}^{\prime}\leftarrow(\mathbf{x}-\mathbf{o})/d$ $\triangleright$ Depth and Ray
6:  $n_{1}\leftarrow n_{1}\perp\mathbf{d}^{\prime},\ \ n_{2}\leftarrow n_{2}\perp\mathbf{d}^{\prime},n_{1}$ $\triangleright$ Basis of plane
7:  $\Pi,\mathbf{J}\leftarrow n_{1},n_{2}$ $\triangleright$ Tangent plane and Transformation
8:  ${\Sigma}^{\prime}\leftarrow\mathbf{J}\mathbf{W}{\Sigma}\mathbf{W}^{\top}\mathbf{J}^{\top}$ $\triangleright$ Projection of covariance
9:  $b_{h},b_{w}\leftarrow{\Sigma}^{\prime}$ $\triangleright$ Range of affected pixels
@@
13:  $\mathbf{N}_{\mathbf{d}}\leftarrow\xi:\{b_{h},b_{w}\}$ $\triangleright$ Depth buffering
14:  $\mathbf{p}:[\bar{\rho},\bar{d},\bar{r}]\leftarrow[0,0,0]$ $\triangleright$ Render Target
@@
17:   $\partial\mathbf{x}\leftarrow d,\epsilon$ $\triangleright$ Back-projection
18:   $\alpha_{i}\leftarrow\alpha_{\xi_{i}}\cdot\mathbb{G}_{i}(\partial\mathbf{x})$ $\triangleright$ Splatting and $\alpha$-weighting
19:   $\bar{\rho},\bar{d},\bar{r}]\leftarrow\alpha_{i}\sum_{j=1}^{i-1}(1-\alpha_{j})[\rho_{\xi_{i}},d_{\xi_{i}},r_{\xi_{i}}]$
@@
22:   $\bar{d},\bar{\rho}\leftarrow 0$ $\triangleright$ Ray-drop
@@
4.3 Dynamic Instances Decomposition
@@
5 Experiments
5.1 Experimental Setting
Real-world Large Scene Datasets. We verified the LiDAR-GS using 12 sequences from Waymo Open Dataset [25] and KITTI-360 Dataset [15], which are public real-world datasets for autonomous driving scenarios. Consistent with previous LiDAR re-simulation methods [26, 29], we utilized 50 continuous frames per sequence, designating every 10th frame for validation and the remaining frames for training. Please refer to Table 1 for the specific sequences chosen.
Table 1: Sequences from Waymo and KITTI-360 for evaluation.
@@ 5.1 Experimental Setting
Baseline and Implementation. For static dataset, we used the official settings to conduct experiments with LiDAR-NeRF [26] and AlignMiF [27]. For NFL [7] which have not yet released their source-code, we stripped the dynamic graph from DyNFL [29] and followed its settings to reproduce the results. Additionally, we maintained the experimental settings as described in LiDAR-NeRF [26] to reproduce the results on PCGen [13] and LiDARsim [20]. For dynamics dataset, we reproduced the results using the same data sequences and official settings as DyNFL [29] and LiDAR4D [34], while also recording the official results of LiDARsim [20] and UniSim [32] from DyNFL [29]. Meanwhile, we also compared our method with contemporaneous works, LiDAR-RT [35] and GS-LiDAR [8], by reproducing their results according to the official source code.
@@
Table 2: Average quantitative results for static scenes from the Waymo Open Dataset and KITTI-360.
@@
5.2 Main Results
Figure 7: For dynamic scenes, qualitative comparison between LiDAR-GS and other methods on Waymo sequence. Points are colorized through their intensity from blue (0) to red (40).
Table 3: Average quantitative results for dynamic scenes.
@@ 5.2 Main Results
Table 4: Average training and inference costs. We reach real-time rendering for LiDAR re-simulation tasks.
[…]
@@
Table 2 and Table 3 present our quantitative results on the public static dataset and dynamic dataset, respectively. As shown, our proposed method LiDAR-GS achieves best quantitative performance among most metrics. Meanwhile, Table 4 demonstrates that our approach significantly reduces training and inference time compared to existing explicit reconstruction, NeRF methods and contemporaneous GS methods. This improvement marks a milestone that LiDAR re-simulation can now achiev the same or even higher frame rate than real-world sensors, typically at 10Hz, making large-scale re-simulation of real-world sequences feasible.
Results on Static Dataset. The quantitative comparison on the Waymo dataset with the static scene is displayed in Table 2, illustrating LiDAR-GS has achieved state-of-the-art in most indicators. As shown in Figure 6, LiDAR-GS is superior to previous methods in modeling scene surfaces and intensity, such as roads, vehicles, and poles. We attribute this superior performance to our well-designed GS reconstruction process and efficient attribute prediction.
Results on Dynamic Dataset. The quantitative comparison on the Waymo dataset with dynamic objects is displayed in Table 3. We can find our proposed method achieves remarkable performance on most evaluation metrics, underscoring the realism of our re-simulated LiDAR scans. Although LiDAR-GS scores lower than LiDAR4D on some intensity metrics (attributed to its two-stage convolutional network optimization), our rendered results, as shown in Figure 7, appear more realistic. And our reconstruction efficiency is significantly higher than that of GS-LiDAR, even though the performance of geometric accuracy is comparable.
Results on Novel View Synthesis. We adjusted the free-viewpoint by lateral shifting (Y) and elevating (Z), or change the number of laser beams (from 64 to 32) for LiDAR re-simulation on novel views as shown in Figure 10, demonstrating our effectiveness of NVS in driving scenarios. And supplementary video shows the more results of the re-simulation application, including vehicle place-in, novel view, beam changing and simulated driving.
5.3 Ablation Results
Table 5: Ablations for the proposed method on Waymo1067.
@@ 5.3 Ablation Results
Ablations on Neural Gaussian Representation. We conducted an ablation experiment to remove the NGR. Table 5 supports our assertions in Section 4.1. The introduction of NGR significantly enhances the capability of scene representation. It can effectively consider the impact of incident direction and distance. As shown in Figure 8, if Gaussian properties are not related to above factors, it cannot identify positions of ray-drop, leading to incorrect optimization of intensity and depth, such as road blind spots.
@@
Figure 9: Ablations on the necessity of adding ray-drop estimation.
Ablations on Laser Beam Splatting. We conducted an ablation experiment to remove the LBS. Specifically, we referred to the camera projection method of the vanilla 3DGS [10] and projected Gaussian onto unified pano projection planes for rasterization. Table 5 demonstrates that LBS is more suitable for LiDAR applications. The improvements provided by LBS are illustrated in Figure 8, confirming our assertions in LABEL:{sec:method:laser-beam-splatting}. Additionally, we implemented LBS using CUDA, resulting in only slight reduction in efficiency. The costs of projections are detailed in Table 4.
Ablations on AABB. The results in Table 5 demonstrate that avioding unnecessary pixel coverage by using AABB for Gaussian projection is beneficial for further convergence.
@@
Discussion on the gap between real and simulation. To gauge the gap between real and simulated scans, we employed detection model, INT [31], on the same trajectory of dynamic data to evaluate the difference in detection results between two point cloud detection results relative to manually annotated GT. Table 6 presents that this difference is negligible in this task, where mAP(mean Average Precision), mF1(mean F1 Score) and mAVE(mean Average Velocity Error) are used as detection metric.
Table 6: Objects detection using simulation data.
@@
6 Conclusion
In this paper, we present the LiDAR Gaussian Splatting method for real-time high-fidelity re-simulation of sensor scans in real-world urban road scenes. LiDAR-GS choose range-view representation to organize laser beams, and innovatively proposes a differentiable laser beam splatting based on micro cross-section plane. Furthermore, LiDAR-GS precisely simulates LiDAR properties using Neural Gaussian Representation and introducing some strategies dedicated to Gaussian laser beam, including AABB on projection constraints, and spatial scale regularization. LiDAR-GS has been verified in multiple dynamic and static scenes, and performs well. In summary, this GS-based rendering LiDAR re-simulation method will accelerate the progress of autonomous driving in data closure.
Limitation and Future Work. Due to the single training perspective and the point density that inversely proportional to the scanning distance, our proposed LiDAR-GS still exhibits degradation in occluded and extrapolated regions at far distances. Throughout our proposed LiDAR-GS pipeline, we also find several points worth further adjustments as a secondary priority: Leverage the ability of generative models, such as diffusion model, to address depth occlusion caused by sparse perspective. And by fusing multi-modal sensor information, such as cameras, model can enhance the understanding of scene features.